# Performance and radiation damage mitigation strategy for silicon photomultipliers on LEO space missions

### v2

@@ 9.1 Annealing
In the laboratory, we study the annealing effect by placing the irradiated devices in a climate chamber at a fixed temperature for a certain period and regularly measuring their IV curve. Then we measure again the current and compare it with the SiPM’s current before irradiation at the given over-voltage of 10 V to determine the recovery efficiency of the annealing process. The time evolution of the 10 V over-voltage current for different storage temperatures is shown in figure 23 for three types of SiPMs. In this figure the y-axis shows the normalized current to the current before irradiation and slightly to the irradiation level. The increase of current is equivalent to the DCR increase due to irradiation. It can be noticed that the current is following an exponential decay with time, whose amplitude, time constant, and offset varying with the storage temperature. The data points in figure 23 are therefore fitted with the following function:
@@
Figure 23: Normalized current (to the current before irradiation at t = 0) measured at 10 V over-voltage vs the time after the SiPM irradiation when they are stored at different temperatures (in the legend). The curves are fitted with an exponential function, for different sensors.
@@
In absence of annealing, irradiation scenario, the solution of this equation is given by the linear increase in DCR as a function of the dose described in equation 6.1. During the annealing process, the solution of the equation is given by the exponential decay model described in equation 9.2. Note that over time, the effectiveness of annealing cycles decreases as the lattice becomes more rigid and less responsive to the heating treatment. Moreover, during the annealing cycle the FPA is still taking irradiation. Taking into account all these factors, the resulting DCR from proton irradiation as a function of time/dose for three different annealing strategies is shown in figure 24.
@@ Acknowledgments
NUSES is funded by the Italian Government (CIPE n. 20/2019), by the Italian Ministry of Economic Development (MISE reg. CC n. 769/2020), by the Italian Space Agency (CDA ASI n. 15/2022), by the European Union NextGenerationEU under the MUR National Innovation Ecosystem grant ECS00000041 - VITALITY - CUP D13C21000430001 and by the Swiss National Foundation (SNF grant n. 178918). This study was carried out also in collaboration with the Space It Up project funded by the Italian Space Agency, ASI, and the Ministry of University and Research, MUR, under contract n. 2024-5-E.0 - CUP n. I53D24000060005.

### v3

@@ 9.1 Annealing
In the laboratory, we study the annealing effect by placing the irradiated devices in a climate chamber at a fixed temperature for a certain period and regularly measuring their IV curve. Then we measure again the current and compare it with the SiPM current before irradiation at the given over-voltage of 10 V to determine the recovery efficiency of the annealing process. The time evolution of the 10 V over-voltage current for different storage temperatures is shown in figure 23 for three types of SiPMs. In this figure, the y-axis shows the normalised current, meaning the dark current relative to that measured during the first measurement after irradiation. It can be noticed that the current is following an exponential decay with time, whose amplitude, time constant, and offset varying with the storage temperature. At time $t_{0}=0$ h, the measurement immediately after irradiation is shown. At time $t_{1}>t_{0}$ in the plot corresponds the first annealing measurement. This takes into account the effect of annealing without temperature to which the sensors were subjected during transport from the irradiation site. The data points in figure 23 are therefore fitted with the following function:
@@
Figure 23: Normalised current measured at 10 V over-voltage vs the time after the SiPM irradiation when they are stored at different temperatures (in the legend). The curves are fitted with an exponential function, for different sensors.
@@
In absence of annealing, irradiation scenario, the solution of this equation is given by the linear increase in DCR as a function of the dose described in equation 6.1. Absent annealing can be defined as the equation being solved at a temperature for which self-annealing is negligible. This is evidenced by the absence of annealing observed in our experimental setup for temperatures below a few degrees. During the annealing process, the solution of the equation is given by the exponential decay model described in equation 9.2. Note that over time, the effectiveness of annealing cycles decreases as the lattice becomes more rigid and less responsive to the heating treatment. Moreover, during the annealing cycle the FPA is still taking irradiation. Taking into account all these factors, the resulting DCR from proton irradiation as a function of time/dose for three different annealing strategies is shown in figure 24.
@@ Acknowledgments
NUSES is funded by the Italian Government (CIPE n. 20/2019), by the Italian Ministry of Economic Development (MISE reg. CC n. 769/2020), by the Italian Space Agency (CDA ASI n. 15/2022), by the European Union NextGenerationEU under the MUR National Innovation Ecosystem grant ECS00000041 - VITALITY - CUP D13C21000430001 and by the Swiss National Foundation (SNF grant n. 178918). The authors would like to thank Officina Stellare S.p.A, in particular Andrea Turella, Andrea Cecconello, Alessandro Scudeler, and Giuseppe Crescenzio, for contributing to the design, manufacturing and assembly of the Terzina Telescope.